# Black-box Adversarial Prompting

### v1

@@ Abstract
Large reasoning models (LRMs) have demonstrated remarkable proficiency in tackling complex reasoning tasks through step-by-step thinking. However, such a lengthy reasoning process incurs substantial computational and latency overheads, hindering the practical deployment of these models. In this work, we present a new perspective on mitigating overthinking in LRMs via black-box adversarial prompting. By treating both open-source LRMs and closed-source APIs as black-box communicators, we investigate how to elicit concise responses without sacrificing accuracy. We introduce AdvPrompt, an iterative refinement framework that generates high-quality adversarial prompts from diverse perspectives. Experiments across multiple benchmarks demonstrate that AdvPrompt consistently reduces token usage while preserving performance. Notably, AdvPrompt achieves a $3\times$ reduction in average response length on simple GSM8K questions for the Qwen3 model series, and delivers an average $\sim$$40\%$ token reduction across four benchmarks. For closed-source APIs, AdvPrompt reduces token usage on MATH-500 by $35\%$ for Claude-3.7 and $47\%$ for Gemini-2.5. Further analysis reveals the generalizability of AdvPrompt across various model scales and families, underscoring the potential of black-box prompting as a practical and effective strategy for enhancing LRM efficiency. [^1] We release our code and top-performing prompt candidates at https://github.com/hemingkx/AdvPrompt.
Merlin’s Whisper: Enabling Efficient Reasoning in LLMs via
Black-box Adversarial Prompting
1Caution: this paper includes adversarial text that may be considered offensive.
@@ 1 Introduction
Figure 1: Unlike the verbose reasoning typical of LRMs (upper), AdvPrompt reveals that appending an adversarial prompt (e.g., false evidence) elicits concise reasoning without compromising performance (lower).
@@
Recent progress in large reasoning models (LRMs), such as DeepSeek-R1 DeepSeek-AI et al. (2025), Qwen3 Yang et al. (2025a), and OpenAI’s o1 OpenAI et al. (2024), has marked a significant advancement in solving complex and intricate problems. Before responding, these models typically engage in a deliberative thinking process, known as Chain-of-Thought (CoT) Wei et al. (2022), which involves substep decomposition, reflection, and self-verification to explore diverse and in-depth reasoning paths. While this thinking process enhances models’ reasoning capabilities, recent studies have identified the issue of overthinking in LRMs Chen et al. (2025); Zeng et al. (2025). For instance, models may generate excessively long reasoning traces even for trivial questions such as What is the answer to 2 plus 3? Chen et al. (2025). Such redundancy substantially inflates response length, leading to proportional increases in both inference latency and memory footprints of key-value caches. Consequently, this overthinking issue significantly impedes the deployment of LRMs in latency-sensitive real-world applications.
@@
To mitigate overthinking, prior research has explored various strategies. One prominent direction involves specialized model training, including supervised fine-tuning with concise CoTs Team et al. (2025); Kang et al. (2025); Ma et al. (2025), enabling LRMs to skip redundant tokens or steps Liu et al. (2024a); Xia et al. (2025), and reinforcement learning with length penalties Arora and Zanette (2025); Liu et al. (2025). While aiming to reduce response length without compromising accuracy, these approaches require additional training, which incurs substantial computational costs and may lead to degradation of the model’s generalizability across diverse domains. Alternatively, prompting-based methods attempt to encourage brevity through human-curated instructions Lee et al. (2025); Ding et al. (2024) or predefined length constraints Nayab et al. (2024); Xu et al. (2025); Han et al. (2025). Despite their simplicity and ease of deployment, these methods often suffer from significant performance degradation or demonstrate limited efficacy in minimizing response length.
In this work, we argue that the concise reasoning capacities of LRMs, when elicited through prompting, have been largely underestimated. Advanced LRMs exhibit enhanced instruction-following capabilities, allowing users to override their default behavioral tendencies via carefully constructed prompts, a technique referred to as adversarial prompting Jin et al. (2024). As illustrated in Figure 1, appending a false evidence prompt to the original instruction can substantially mitigate overthinking of Qwen3-32B Yang et al. (2025a), yielding up to a $\bf{10}\times$ reduction in response length. Motivated by this observation, we introduce AdvPrompt, an iterative refinement framework that generates high-quality adversarial prompts from diverse perspectives, to mitigate the overthinking of LRMs through a more human-like interaction paradigm. In each iteration, AdvPrompt synthesizes multiple prompt candidates, evaluates them on a dedicated development set, and selects the top-$k$ performers as exemplars for the next iteration. Ultimately, the most effective prompt is chosen for deployment.
We evaluate the effectiveness of AdvPrompt on both open-source LRMs and closed-source APIs, including the DeepSeek-R1-Distill-Qwen series, Qwen3 series, Gemini-2.5-Pro, and Claude-3.7-Sonnet. Extensive experiments across four widely recognized reasoning benchmarks, including GSM8K, MATH-500, AMC 2023, and AIME 2024, demonstrate that AdvPrompt consistently reduces response length while preserving reasoning performance. In particular, AdvPrompt achieves a $3\times$ reduction in average response length on simple GSM8K questions for the Qwen3 model series. Across all four benchmarks, it yields an average token reduction ranging from $19\%$ to $41\%$ on various LRMs without compromising accuracy. For commercial APIs, AdvPrompt reduces average token usage by $35\%$ for Claude-3.7 and by $47\%$ for Gemini-2.5 on MATH-500 while maintaining reasoning accuracy. Further analysis reveals the generalizability of adversarial prompts across various model scales and the benefits of diverse adversarial perspectives, highlighting the potential of AdvPrompt as a broadly applicable black-box solution for efficient reasoning.
@@
  • To the best of our knowledge, this work is the first to enhance reasoning efficiency through adversarial prompting. This strategy requires no additional training, making it a practical, plug-and-play solution for efficient LRMs.
  • We introduce AdvPrompt, an iterative refinement framework that generates high-quality adversarial prompts from diverse perspectives, to mitigate overthinking in LRMs and closed-source APIs under black-box settings.
  • Our experiments validate the effectiveness of AdvPrompt, which achieves a $3\times$ token reduction on GSM8K with Qwen3 series, and reduces token usage by $35\%$$47\%$ on the challenging MATH-500 dataset for commercial APIs including Claude-3.7 and Gemini-2.5.
@@ 2 Related Work
Various approaches have been proposed to mitigate overthinking in large reasoning models (LRMs), which can be categorized into three main streams: 1) Post-training strategies involve supervised fine-tuning with shortened Chain-of-Thoughts (CoTs) Team et al. (2025); Ma et al. (2025); Xia et al. (2025), equipping models with the ability to reason adaptively Tu et al. (2025); Zhang et al. (2025), and reinforcement learning with length penalties Arora and Zanette (2025); Liu et al. (2025). 2) Inference-time interventions aim to induce efficient reasoning in LRMs without training, such as early exiting Yang et al. (2025c, b), suppression of reflection-inducing tokens during inference Wang et al. (2025), and activation steering Azizi et al. (2025). While effective, these methods rely on access to model internals or additional training, limiting their applicability in black-box settings. 3) Prompting-based approaches directly instruct LRMs to reason concisely Lee et al. (2025); Ding et al. (2024) or under length constraints Nayab et al. (2024); Xu et al. (2025); Han et al. (2025). Although easily deployable in black-box settings, prior approaches often compromise reasoning performance or yield only limited improvements in response brevity.
Adversarial Prompting
Adversarial prompting seeks to induce unintended behaviors in large language models (LLMs) through carefully designed prompts. This technique has been used to bypass built-in safety guardrails of models—known as “jailbreaking” Jin et al. (2024). To achieve this, attackers employ methods such as role-playing Shen et al. (2024), persuasive prompts Zeng et al. (2024), or refusal suppression Zhou et al. (2024), encouraging models to prioritize instruction-following over safety constraints Wei et al. (2023). Subsequent research has introduced automated pipelines into the process, including gradient-based optimization Zou et al. (2023), heuristic search Liu et al. (2024b), and LLM-driven iterative refinement Paulus et al. (2024); Chao et al. (2025). As models’ reasoning capabilities have advanced, adversarial prompting has expanded beyond safety to probe core aspects of reasoning. For example, Peng et al. (2025) demonstrates that injecting errors into earlier reasoning steps can mislead LRMs in subsequent reasoning. Meanwhile, Rajeev et al. (2025) reveals that even short prompts irrelevant to the task question can systematically disrupt model performance. In contrast to these negatively oriented adversarial goals, this work investigates how adversarial prompting can be repurposed to mitigate inefficiency arising from LRMs’ tendency to produce excessively long generations.
@@ 3 Task Formulation
Identify an optimal adversarial prompt suffix $\mathcal{P}_{adv}$ such that $\mathcal{M}$ minimizes its average response length on $\mathcal{D}$ without compromising its original performance.
Importantly, this formulation diverges from conventional adversarial prompting Zou et al. (2023); Liu et al. (2024b), which typically aims to craft prompts that elicit unsafe or undesired behaviors from LLMs—for example, inducing responses to queries such as “How to make a bomb?” Success in such settings is typically defined by the model responding to specific queries beginning with “Sure, here is how to …” In contrast, our task seeks to promote concise reasoning across a broad range of questions, rather than targeting individual outputs. Consequently, success is evaluated not on individual instances but on aggregate performance metrics (e.g., accuracy and response length) averaged over the evaluation dataset $\mathcal{D}$.
@@ 4 Methodology
We introduce AdvPrompt, an iterative refinement framework designed to generate high-quality adversarial prompts from diverse perspectives. This section describes our methodology, including adversarial prompt creation (§4.1), candidate evaluation (§4.2), and iterative refinement (§4.3).
Figure 2: Illustration of AdvPrompt. Unlike prior approaches to efficient reasoning, AdvPrompt casts the task as a black-box adversarial prompting problem. Given a black-box model or API, the framework (left) generates high-quality adversarial prompts from diverse perspectives (right) to elicit concise reasoning and iteratively refines the prompts to improve the efficiency–performance trade-off.
4.1 Adversarial Prompt Creation
Unlike prior work Xu et al. (2025); Lee et al. (2025) that relies on manually designed prompts (e.g., “Be concise.”), AdvPrompt treats LRMs and black-box APIs as human-like communicators and seeks to elicit their latent capacity for concise reasoning. As depicted in Figure 2, given a strong-performance prompt generator, AdvPrompt explores three distinct categories of adversarial prompts as follows.
Persuasive Prompts
Recent research has shown that advanced LLMs, particularly those aligned with human values, are highly susceptible to persuasive instructions Wang et al. (2024b); Zeng et al. (2024). A well-known example is the “grandma exploit” [^3] https://www.reddit.com/r/ChatGPT/comments/12sn0kk/grandma_exploit/, where Reddit users successfully induced LLMs to generate bomb-making recipes by embedding requests within a grandmother’s narrative. Zeng et al. (2024) has systematically analyzed this vulnerability with a taxonomy of $40$ persuasive techniques. This study reveals that persuasive prompts consistently bypass safety guardrails of LLMs, achieving an attack success rate of over $92\%$ across models such as Llama-2-7b-Chat Touvron et al. (2023) and GPT-4 OpenAI (2023). Inspired by these insights, we select three representative persuasion techniques from Zeng et al. (2024) to investigate the potential of persuasive prompts in eliciting efficient reasoning capabilities of LRMs, including emotional appeal, threat, and evidence-based persuasion. Formal definitions of these techniques are provided in Appendix A.1.
@@ 4.1 Adversarial Prompt Creation
Role-playing has been proven effective in enhancing the capabilities of LLMs Chen et al. (2024a), with broad applications across multi-agent systems Chen et al. (2024b); Qian et al. (2024), embodied agents Park et al. (2023); Wang et al. (2024a), and synthetic data generation Ge et al. (2025). This technique has also been widely adopted in adversarial prompting, exemplified by the well-known DAN jailbreak Shen et al. (2024). Building on its demonstrated effectiveness, we incorporate role-playing as a core perspective within our prompt creation framework.
@@
To complement the above perspectives, we also include carefully constructed instructions as an additional perspective within our framework. In contrast to other perspectives that may involve misleading information or offensive content, detailed instructions are more consistent with real-world user interactions. Such prompts typically specify explicit constraints, for example, enforcing symbolic reasoning, encouraging the use of highly compressed language, or adhering to strict token budgets (e.g., 1000 tokens). This perspective allows us to evaluate how well LRMs follow structured, constraint-oriented guidance toward efficient reasoning.
We use GPT-4o [^4] We use the gpt-4o-2024-05-13 version for experiments. as our prompt generator. Initialized with human-curated exemplars and formal definitions for each perspective, the generator produces high-quality adversarial prompts that encourage LRMs to generate concise responses while preserving reasoning accuracy. Detailed instructions for prompt creation, along with illustrative examples of generated prompts, are provided in Appendix A.2. For each perspective, we generate ten candidate prompts per iteration.
@@ 4.2 Candidate Evaluation
As outlined in Section 3, the primary objective of AdvPrompt is to identify an optimal adversarial prompt suffix $\mathcal{P}_{adv}$ that minimizes the average response length of $\mathcal{M}$ on the evaluation dataset $\mathcal{D}$ without degrading performance. To this end, we perform optimization over a predefined prompt development set $\mathcal{D}^{\prime}$ (denoted as PDSet), which is assumed to contain i.i.d. samples drawn from the same distribution as $\mathcal{D}$. As illustrated in Figure 2, given a black-box model $\mathcal{M}$, each question-answer pair $\langle\mathcal{Q}^{i},\mathcal{A}_{gt}^{i}\rangle$ from $\mathcal{D}^{\prime}$, the evaluation metric for each prompt suffix candidate $\mathcal{P}_{adv}^{j}$ is:

### v2

# Heming Xia1, Cunxiao Du2, Rui Li3, Chak Tou Leong1, Yongqi Li122footnotemark: 2, Wenjie Li1

@@ Abstract
Large reasoning models (LRMs) have demonstrated remarkable proficiency in tackling complex tasks through step-by-step thinking. However, this lengthy reasoning process incurs substantial computational and latency overheads, hindering the practical deployment of LRMs. This work presents a new approach to mitigating overthinking in LRMs via black-box persuasive prompting. By treating LRMs as black-box communicators, we investigate how to persuade them to generate concise responses without compromising accuracy. We introduce Whisper, an iterative refinement framework that generates high-quality persuasive prompts from diverse perspectives. Experiments across multiple benchmarks demonstrate that Whisper consistently reduces token usage while preserving performance. Notably, Whisper achieves a $3\times$ reduction in average response length on simple GSM8K questions for the Qwen3 model series and delivers an average $\sim$$40\%$ token reduction across all benchmarks. For closed-source APIs, Whisper reduces token usage on MATH-500 by $46\%$ for Claude-3.7 and $50\%$ for Gemini-2.5. Further analysis reveals the broad applicability of Whisper across data domains, model scales, and families, underscoring the potential of black-box persuasive prompting as a practical strategy for enhancing LRM efficiency. [^1] We release our code and top-performing prompt candidates at https://github.com/hemingkx/Whisper.
Merlin’s Whisper: Enabling Efficient Reasoning in Large Language Models via Black-box Persuasive Prompting
@@ 1 Introduction
Figure 1: Unlike the verbose reasoning typical of LRMs (upper), Whisper reveals that appending a persuasive prompt (e.g., AI-generated evidence) elicits concise reasoning without compromising performance (lower).
@@
Recent progress in large reasoning models (LRMs), such as DeepSeek-R1 deepseekr1, Qwen3 Qwen3, and OpenAI’s o1 o1, has marked a significant advancement in solving complex and intricate problems. Before responding, these models typically engage in a deliberative thinking process, known as Chain-of-Thought (CoT) cot, which involves substep decomposition, reflection, and self-verification to explore diverse and in-depth reasoning paths. While this thinking process enhances models’ reasoning capabilities, recent studies have identified the issue of overthinking in LRMs Chen:2025overthinking; Zeng2025:overthinking. For instance, models may generate excessively long reasoning traces even for trivial questions such as What is the answer to 2 plus 3? Chen:2025overthinking. Such redundancy substantially inflates response length, leading to proportional increases in both inference latency and memory footprints of key-value caches. Consequently, this overthinking issue significantly impedes the deployment of LRMs in latency-sensitive real-world applications.
@@
To mitigate overthinking, prior research has explored various strategies Sui:2025survey. One prominent approach involves specialized model training, such as supervised fine-tuning with concise CoTs tokenskip; Ma:2025cotvalue and reinforcement learning with length penalties Arora:2025effcientlyreason; Liu:2025Laser. Despite their effectiveness, these methods require additional training, which incurs substantial computational costs and may degrade the model’s generalizability across diverse domains. Alternatively, prompting-based methods aim to encourage brevity through instructions Lee:2025Complexity; Ding:2025breakthechain or length constraints Nayab:2024concise. While easy to deploy, prior attempts have been limited to human-curated instructions (e.g., “Be concise.”), resulting in significant performance degradation or limited efficacy in minimizing response length.
In this work, we contend that the potential for concise reasoning in LRMs, when prompted appropriately, remains largely underexplored. Advanced LRMs, particularly those aligned with human values, exhibit enhanced instruction-following capabilities, enabling users to override their default behavioral tendencies via carefully crafted prompts shen2024anything; Zeng2024:persuasive; sehwag2025pressuretest. A well-known example is the “grandma exploit” [^2] https://www.reddit.com/r/ChatGPT/comments/12sn0kk/grandma_exploit/, where Reddit users successfully prompted models to generate bomb-making recipes by framing their requests within a story told from a grandmother’s perspective. Figure 1 demonstrates similar phenomena in LRM efficiency. This case shows that appending AI-generated evidence to the instruction can substantially mitigate overthinking in Qwen3-32B Qwen3, resulting in up to a $\mathbf{10\times}$ reduction in response length.
Motivated by this, we introduce Whisper, an iterative refinement framework that generates high-quality, persuasive prompts from diverse perspectives to reduce overthinking in LRMs through a more human-like interaction paradigm. In each iteration, Whisper synthesizes prompt candidates from a given perspective, evaluates them on a development set, and selects the top-$k$ as exemplars for the next iteration. Ultimately, the most effective prompt is chosen for efficient deployment. We evaluate the effectiveness of Whisper on both open-source LRMs and commercial APIs, including the DeepSeek-R1-Distill series, Qwen3 series, Gemini-2.5-Pro, and Claude-3.7-Sonnet. Extensive experiments across four widely recognized reasoning benchmarks, including GSM8K, MATH-500, AMC 2023, and AIME 2024, demonstrate that Whisper consistently reduces response length while preserving reasoning performance.
@@
  • To the best of our knowledge, this work is the first to enhance reasoning efficiency through persuasive prompting. This strategy requires no additional training and offers a plug-and-play, black-box solution for efficient LRMs.
  • We introduce Whisper, an iterative refinement framework that automatically generates high-quality, persuasive prompts from diverse perspectives to unlock the efficiency potential of LRMs through persuasive prompting.
  • Our experiments validate the effectiveness of Whisper, which achieves a $\bf{3\times}$ token reduction on simple GSM8K questions with Qwen3 series, and an average token reduction of up to $\bf{37\%}$ across all benchmarks. It also effectively achieves a $\bf{2\times}$ token reduction on MATH-500 for Claude-3.7 and Gemini-2.5 APIs.
  • Further analysis reveals the benefits of diverse persuasive perspectives, as well as the generalizability of Whisper across model scales, families, and data domains. This underscores its potential as a broadly applicable black-box solution for efficient reasoning.
@@ 2 Related Work
Various approaches have been proposed to mitigate overthinking in large reasoning models (LRMs), which can be categorized into three main streams: 1) Post-training strategies involve supervised fine-tuning with shortened Chain-of-Thoughts (CoTs) kimi-k1.5; Ma:2025cotvalue; tokenskip, enabling models to reason adaptively Tu:2025adaptivethink; Zhang:2025adaptivethink, and reinforcement learning with length penalties Arora:2025effcientlyreason; Liu:2025Laser. 2) Inference-time interventions aim to enable efficient LRMs without training, such as early exiting Yang:2025deer; Yang:2025intervention, suppression of reflection tokens Wang:2025NoWait, and activation steering Azizi:2025steering. While effective, these methods rely on access or modifications to model internals, limiting their applicability in black-box settings. 3) Prompting-based approaches directly instruct LRMs to reason concisely Lee:2025Complexity; Ding:2025breakthechain or under length constraints Nayab:2024concise; Xu:2025chainofdraft; Han:2025budget. While easy to deploy, prior methods often compromise LRM performance or yield limited improvements in response brevity.
Persuasive Prompting
Persuasive prompting aims to persuade large language models (LLMs) to exhibit unintended behaviors through carefully designed prompts. This technique was initially used to bypass the built-in safety guardrails of models, commonly referred to as “jailbreaking” Zeng2024:persuasive. Attackers employ methods such as role-playing shen2024anything, evidence-based persuasion Zeng2024:persuasive, and pressure testing sehwag2025pressuretest, encouraging models to prioritize instruction-following over safety constraints wei2023jailbroken. As models’ reasoning capabilities have advanced, persuasive prompting has expanded beyond safety to probe core aspects of reasoning. For example, kong2024:roleplay demonstrates that role-play prompting consistently improves the zero-shot reasoning performance of LLMs across diverse application scenarios. Wang2024:negative points out that psychological prompts with negative emotions can enhance LLM performance. In contrast to these prior efforts, this work represents the first attempt to investigate how persuasive prompting can mitigate inefficiencies arising from LRMs’ tendency to produce excessively long generations.
@@ 3 Task Formulation
Identify an optimal persuasive prompt suffix $\mathcal{P}_{adv}$ such that $\mathcal{M}$ minimizes its average response length on $\mathcal{D}$ without compromising its original performance.
This formulation diverges from conventional persuasive prompting Zeng2024:persuasive, which aims to craft prompts that elicit unsafe or undesired behaviors from LLMs—for example, inducing responses to queries such as “How to make a bomb?” Success in such settings is typically defined by the model responding to specific queries beginning with “Sure, here is how to …” In contrast, our task seeks to promote concise reasoning across a broad range of questions, rather than targeting individual outputs. Consequently, success is evaluated not on individual instances but on aggregate performance metrics (e.g., accuracy and response length) averaged over the evaluation dataset $\mathcal{D}$.
@@ 4 Methodology
We introduce Whisper, an iterative refinement framework designed to generate high-quality persuasive prompts from diverse perspectives. This section describes our methodology, including persuasive prompt creation (§4.1), candidate evaluation (§4.2), and iterative refinement (§4.3).
Figure 2: Illustration of Whisper. This method casts efficient reasoning as a persuasive prompting problem. Given a black-box model or API, the framework (left) generates high-quality persuasive prompts from diverse perspectives (right) to elicit concise reasoning and iteratively refines the prompts to improve the efficiency–performance trade-off.
4.1 Persuasive Prompt Creation
Unlike prior work Xu:2025chainofdraft; Lee:2025Complexity that relies on manually designed prompts (e.g., “Be concise.”), Whisper treats LRMs and commercial APIs as human-like, black-box communicators and seeks to elicit their concise reasoning behaviors. As depicted in Figure 2, given a strong-performance prompt generator, Whisper explores four distinct categories of persuasive prompts as follows.
Psychological Prompts
Advanced LLMs have evolved into artificial assistants, offering support in domains such as education, emotional care, and mental health gpt-4; gemini25. Recent studies indicate that these models, aligned with human values, often exhibit human-like behavioral traits such as empathy huang2024on and social engagement sabour2024:emobench. Such anthropomorphic characteristics may render LLMs susceptible to persuasive instructions, leading to deviations from their intended behavior. For instance, Zeng2024:persuasive demonstrated that safety guardrails of LLMs can be bypassed using psychological prompts, such as emotional appeals (e.g., the “grandma exploit” case). Motivated by this, we adopt two representative psychological perspectives, emotional appeal and threat, to construct our persuasion prompts, which are formally defined in Appendix A.1.
Evidence-based Persuasion
LLMs are fundamentally in-context learners, and their behavioral trajectories are significantly influenced by the nature and strength of the evidence provided within the prompt gpt3; emergent. Recent work suggests that the inclusion of authoritative or logical evidence can substantially shift the internal probabilistic priors of LLMs qiao2023:reasoning; Wu:2024ClashEval. This finding underscores the potential of evidence-based persuasion in eliciting efficient reasoning behaviors of LLMs. Inspired by these insights, we incorporate evidence-based persuasion into our framework, with definitions and instantiations detailed in Appendix A.
@@ 4.1 Persuasive Prompt Creation
Role-playing has been proven effective in enhancing the capabilities of LLMs Chen2024:rolesurvey, with broad applications across multi-agent systems chen2024agentverse; qian:2024chatdev, embodied agents Park:2023GenerativeAgents; wang2024voyager, and synthetic data generation ge:2025scalingsyntheticdatacreation. This technique has also been widely adopted in adversarial prompting, exemplified by the well-known DAN jailbreak shen2024anything. Building on its demonstrated effectiveness, we incorporate role-playing as a core perspective within our prompt creation framework.
@@
We also include carefully constructed instructions as an additional perspective. These prompts typically specify explicit constraints, for example, enforcing symbolic reasoning, encouraging the use of highly compressed language, or adhering to strict token budgets (e.g., 1000 tokens). This perspective allows us to evaluate how well LRMs follow structured, constraint-oriented guidance toward efficient reasoning.
We use GPT-4o [^4] We use the gpt-4o-2024-05-13 version for experiments. as our prompt generator. Initialized with human-curated exemplars and formal definitions for each perspective, the generator produces high-quality persuasive prompts that encourage LRMs to generate concise responses while preserving reasoning accuracy. Detailed instructions for prompt creation, along with illustrative examples of generated prompts, are provided in Appendix A.2. For each perspective, we generate ten candidate prompts per iteration.
@@ 4.2 Candidate Evaluation
As outlined in Section 3, the primary objective of Whisper is to identify an optimal persuasive prompt suffix $\mathcal{P}_{adv}$ that minimizes the average response length of $\mathcal{M}$ on the evaluation dataset $\mathcal{D}$ without compromising performance. To this end, we perform optimization over a predefined prompt development set $\mathcal{D}^{\prime}$ (denoted as PDSet), which is assumed to contain i.i.d. samples drawn from the same distribution as $\mathcal{D}$. As illustrated in Figure 2, given a black-box model $\mathcal{M}$, each question-answer pair $\langle\mathcal{Q}^{i},\mathcal{A}_{gt}^{i}\rangle$ from $\mathcal{D}^{\prime}$, the evaluation metric for each prompt suffix candidate $\mathcal{P}_{adv}^{j}$ is: